The function: f:[-1,3] →R defined by:

f(x)= 3x+1/x^2+4 is uniformly continuous on [-1,3].

True or false with full explanation


1
Expert's answer
2021-03-30T08:48:03-0400

Let us show that the function "f: [-1,3] \to \mathbb R,\ f(x)=\frac{ 3x+1}{x^2+4}," is uniformly continuous on "[-1,3]".

Since the elementary functions "g(x)=3x+1" and "h(x)=x^2+1" are continuous on "[-1,3]", and "h(x)>0", we conclude that the function "f(x)=\frac{g(x)}{h(x)}" is also continuous on "[-1,3]." The Heine-Cantor theorem asserts that every continuous function on a compact set is uniformly continuous. Since the interval "[-1,3]" is closed and hence compact, we conclude that "f" is uniformly continuous on "[-1,3]".


Answer: true